QUESTION 1                                                                                                                      

AA Entity and all its suppliers are registered VAT vendors. Where applicable, VAT is charged at 15%.

AA Entity sells crockery at a mark-up percentage of 25% gross profit on cost price. AA Entity is insured for potential inventory losses due to fire for an amount of R218 500 (including VAT).

On 21 July 2019, a fire broke out in the store room of AA Entity which destroyed almost all of AA Entity’s trade inventories, except for trade inventories with a cost price of R80 000 that the entity could salvage. This salvaged trade inventory items are still in a very good condition and will be sold at the normal selling price.

AA Entity uses the periodic inventory system and therefore does not know what the value of trade inventories on hand was on 21 July 2019. The last inventory count was performed at the reporting date on 30 June 2019 and the inventory records reflected trade inventories on hand at a cost price of R341 000 on that date.

The total value of supplier’s invoices for trade inventories purchased since the reporting date until the date of the fire, amounted to R59 280. The total value of sales invoices made out to customers since the reporting date until the date of the fire amounted to R159 600.

 

REQUIRED

1.1Calculate the value of the insurance compensation that will probably be paid by the insurance company in respect of the trade inventories destroyed in the fire.    



1
Expert's answer
2021-01-05T14:27:00-0500

cost of inventory after fire:

341 000-80 000=261 000

stocks are subject to VAT and mark-ups:

"261 0000\times0.25+261000=326250"

"326250\times0,15+326250=375187.50"

VAT deductible:

159600-59280=100320

amount of damage:

375187.50-100320=274867.50

insurance compensation under the proportional liability system:

"Q=\frac{T\times S}{W}"

T=274 867.50

S=218 500

W=375 187.50

"Q=\frac{274 867.50\times 218 500}{375187.50}=160 076.09"